# Resource Management and Circuit Scheduling for Distributed Quantum Computing Interconnect Networks

### v3

@@ V Single-QCirc Assignment
Figure 2: MILP-based Batch-QCirc assignment problem formulation.
{mdframed}
[linewidth=0.5pt, roundcorner=5pt]
$\min~{}(\frac{1}{T_{\rm dec}})\sum_{j_{1}}{\sum_{j_{2}>j_{1}}{z_{mj_{1}j_{2}}T_{j_{1}j_{2}}}}$
s.t.
$\sum_{j_{1}}{\sum_{j_{2}>j_{1}}{z_{mj_{1}j_{2}}(1-F_{j_{1}j_{2}})}}\leq th$
$\sum_{j=1}^{J}{x_{mj}}=w_{m}$
$x_{mj}\leq n_{j}\quad\text{for}\quad j=1,{\scriptstyle\ldots},J$
$\sum_{j=1}^{J}{r_{mj}}\leq K_{m}^{\rm max}$
$r_{mj}\leq x_{mj}\leq\gamma r_{mj}\;\;\text{for}\;\;j=1,{\scriptstyle\ldots},J$
$z_{mj_{1}j_{2}}\leq r_{mj_{1}}\;\text{for}\;\;j_{1},j_{2}=1,{\scriptstyle\ldots},J\;j_{2}>j_{1}$
$z_{mj_{1}j_{2}}\leq r_{mj_{2}}\;\text{for}\;\;j_{1},j_{2}=1,{\scriptstyle\ldots},J\;j_{2}>j_{1}$
$z_{mj_{1}j_{2}}\geq r_{mj_{1}}+r_{mj_{2}}-1\;\text{for}\;j_{1},j_{2}=1,{\scriptstyle\ldots},J\;j_{2}>j_{1}$
@@
Figure 3: MILP-based Single-QCirc assignment problem formulation.
VI Evaluation and results
@@
This section evaluates the proposed scheduling algorithms through simulations. We consider a DQC network consisting of 16 QPUs within a data center. The QPUs have the following capacities: four QPUs with capacity 8, four with capacity 12, four with capacity 16, and four with capacity 20. These capacities are randomly assigned to QPUs using a fixed random seed to ensure consistent QPU configurations across all simulations. The entire simulation process is repeated for 10 different seed values, and the average results are presented.
@@
The network adopts a 4-pod fat-tree topology [28], as illustrated in Fig. 5. In Sec. II, we defined two key parameters characterising the links in the logical topology: latency ($T_{j_{1}j_{2}}$) and fidelity ($F_{j_{1}j_{2}}$). Here, we specify their values for the 4-pod fat-tree topology used in our simulations. This topology features three types of links, distinguished by the number of optical switches along the path—specifically, 1, 3, or 5 switches. We denote the corresponding latency values by $T_{\text{link}_{1}}$, $T_{\text{link}_{3}}$, and $T_{\text{link}_{5}}$, respectively. Assuming an elementary link latency of $T_{\text{el}}$ (corresponding to a link without any optical switches), and noting that the entanglement generation rate is directly affected by link loss, the latency for each link type is calculated using the formula:
$T_{\text{link}_{n_{s}}}=\frac{T_{\text{el}}}{\eta_{s}^{n_{s}}},$
@@
where $\eta_{s}$ is the switch transmission efficiency and $n_{s}\in\{1,3,5\}$ is the number of switches in the path. In our simulations, we assume that all optical switches are of the same type and have an identical loss of $0.5$ dB. Given the short physical distances between QPUs in a data center environment, propagation delays are considered negligible. For the fidelity parameter, we assume an identical value across all links.
To establish the set of quantum circuits, $\mathcal{QC}$, benchmark circuits from the Munich toolkit [29] are used. The simulation considers a pool of four distinct QCirc types: Quantum Fourier Transform (QFT), Deutsch-Jozsa (DJ), WState, and GHZ state. QCircs are randomly selected from this pool, with the number of qubits $w_{m}$ for each circuit sampled from a range $\mathcal{R}_{w}$. Two cases are studied: (a) $\mathcal{R}_{w}=[20,30]$ and $\mathcal{R}_{w}=[30,40]$. In both cases, the value of $w_{m}$ is reduced by 10 whenever a QFT circuit is selected—that is, a reduced range is effectively used for QFT—to account for the faster scaling of $\nu_{m}$ with $w_{m}$. As before, simulations are repeated across 10 fixed random seeds, and the results are averaged.
To solve the MILP optimisation problems presented in Problem Formulations 2 and 3, the Python-MIP package with the Gurobi optimiser is used. The threshold for $C_{\rm asg}$, denoted as $\kappa$, is initially set to $M$, and it is decremented until a valid solution is found. Once the optimal assignment is obtained, we proceed to partition each quantum circuit individually using the Kernighan-Lin algorithm provided by the NetworkX library. This algorithm determines both the number of remote gates between QPU pairs and the total number of remote gates for each QCirc. For circuits requiring partitioning into more than two parts, the Kernighan-Lin algorithm is applied iteratively until the desired number of partitions is achieved.
@@
To evaluate the performance of the proposed scheduling methods, we consider several key figures of merit. All results are averaged over repeated simulation runs.
@@
  • Normalised Job Execution Time (JET) for each QCirc type: Average execution time for each QCirc type, normalised by $T_{\rm dec}$, i.e., JET/$T_{\rm dec}$.
@@
  • Normalised Throughput: defined as the ratio of the number of QCircs, $M$, to Normalised Makespan.
  • QPU usage, defined as follows:
    QPU $\text{QPU Usage}=\frac{\sum_{j}{N_{j}\times\text{TBT}(\text{QPU}_{j})}}{(\sum_{j}{N_{j}})(\text{Makespan})}$ (6)
Here, “TBT” is Total Busy Time for a QPU. As a baseline, random scheduling is used for comparison.
The normalised JET for a specific QCirc is obtained under the following assumptions. First, we assume that remote gates within a layer of a given QCirc are executed sequentially rather than in parallel, in accordance with current and near-future technological limitations. Second, we assume that all local gates have identical execution durations. Based on these assumptions, we iterate through all circuit layers and identify whether each layer contains any remote gates or only local gates. Let $N_{\rm LL}$ denote the number of layers containing only local gates. We compute the normalised JET using the following expression:
@@
Figure 6 presents the average remote gate count and partition count per QCirc for the proposed scheduling methods, compared against random scheduling as a baseline. For the batch-QCirc assignment method, several values of the parameter $\alpha$ are considered. As shown, both the batch-QCirc and single-QCirc methods yield significant improvements in these metrics. Moreover, in all cases, the batch-QCirc method outperforms the single-QCirc approach. These results are particularly important, as the number of remote gates directly reflects the demand for communication resources. Therefore, the findings demonstrate that both proposed methods—and especially the batch-QCirc assignment—effectively reduce communication overhead.
Figure 6: Average number of remote gates and circuit partitions per QCirc for the proposed and baseline scheduling methods. For the batch-QCirc approach, several values of the parameter $\alpha$ are considered. (a)–(b): $\mathcal{R}_{w}=[20,30]$, (c)–(d): $\mathcal{R}_{w}=[30,40]$.
Figure 7: Average Job Execution Time (JET), normalised by $T_{\text{dec}}$, for the proposed and baseline scheduling methods across different circuit types. (a)–(d): $\mathcal{R}_{w}=[20,30]$; (e)–(h): $\mathcal{R}_{w}=[30,40]$.
@@
Figure 8: Average Makespan, Throughput, and QPU Usage for the proposed and baseline scheduling methods. Makespan and, by extension, Throughput are normalised by $T_{\text{dec}}$. (a)–(c): $\mathcal{R}_{w}=[20,30]$; (d)–(f): $\mathcal{R}_{w}=[30,40]$.
Next, we evaluate the normalised Job Execution Time (JET) across all circuit types. As shown in Fig. 7, the proposed methods generally outperform the baseline. Specifically, for the qubit range $\mathcal{R}_{w}=[20,30]$, the JET is reduced in almost all cases, with batch-QCirc scheduling outperforming the single-QCirc approach.
For $\mathcal{R}_{w}=[30,40]$, the results vary by circuit type. For QFT and DJ circuits, batch-QCirc scheduling achieves the lowest JET, while single-QCirc still performs better than random scheduling. In contrast, for W-state and GHZ circuits, the best-performing method depends on both the circuit type and the value of $M$. For example, in the case of GHZ circuits, the single-QCirc scheduler achieves the lowest JET in nearly all cases, except when $M=5$.
It is also worth noting that the normalised JET values for QFT circuits are higher overall compared to GHZ circuits, which aligns with their corresponding $\nu_{m}$ values. These results highlight the system-level, holistic nature of resource management in batch-QCirc scheduling, where trade-offs may arise—for instance, reducing the JET of highly connected circuits at the expense of not fully optimising the JET of circuits with lower connectivity.
Overall, these results demonstrate the effectiveness of the proposed methods in reducing the JET of quantum circuits. This improvement is particularly significant for circuits with higher $\nu_{m}$ values, such as the QFT circuits in our simulations. It can be observed that under random scheduling, the normalised JET (i.e., JET/$T_{\text{dec}}$) often exceeds 1, which may render successful execution infeasible due to decoherence. However, by applying the proposed methods, this ratio is significantly reduced, enabling more reliable circuit execution within the coherence time constraints.
Next, we examine the normalised makespan, normalised throughput, and QPU usage. The results are shown in Fig. 8. Compared to random scheduling, the proposed methods significantly improve all three metrics in nearly all cases. Additionally, the batch-QCirc approach outperforms the single-QCirc variant in most scenarios, especially when $\mathcal{R}=[20,30]$. These improvements are primarily driven by the reduction in JET.
It is important to note that QPU usage alone is not a sufficient indicator of resource efficiency; it should be interpreted alongside the makespan to enable a proper evaluation of scheduling performance.
@@
Lastly, we compare the results for batch-QCirc scheduling with varying values of $\alpha$, as presented across Figures 6, 7, and 8. In terms of makespan, throughput, and QPU usage, $\alpha=0.75$ consistently yields the worst performance, whereas $\alpha=0.55$ or $\alpha=0.65$ typically perform best. Since the parameter $\alpha$ determines when the algorithm proceeds to the next scheduling cycle, it plays a crucial role in these metrics. Looking at the results for remote gate count and JET, $\alpha=0.75$ does not show meaningful improvements over $\alpha=0.55$ or $\alpha=0.65$; in fact, the latter two often outperform it. These findings indicate that increasing $\alpha$ too much can lead to unnecessary resource waste without performance gains.
VII Conclusion
In this work, we addressed the challenge of efficient resource management and QCirc scheduling in distributed quantum computing (DQC) networks, with a particular focus on computing resources. We proposed two scheduling approaches: a dynamic batch-QCirc scheduling method, which operates through sequential scheduling cycles with batch optimisation, and a single-QCirc scheduling method that optimises QCirc allocation individually.

### v4

@@ III-A1 Batch-QCirc Assignment
Formulation 1. MILP-based Batch-QCirc assignment problem $\min~\bigg{\{}\sum_{m\in\mathcal{I}}\sum_{k=1}^{K_{m}^{\rm max}}\sum_{1\leq j_{1}<j_{2}\leq J}\nu_{mk}\,x_{mkj_{1}j_{2}}\big{(}\omega_{0}\frac{w_{m}T_{j_{1}j_{2}}}{T_{\rm dec}}+$ $\qquad\qquad\omega_{1}(1-F_{j_{1}j_{2}})\big{)}\bigg{\}}$ s.t. $(1)~\sum_{m\in\mathcal{I}}\sum_{k=1}^{K_{m}^{\rm max}}y_{mk}=\zeta$ $(2)~\sum_{j=1}^{J}r_{mj}s_{j}N_{j}\;\geq\;w_{m}\sum_{k=1}^{K_{m}^{\rm max}}y_{mk}\quad\forall m\in\mathcal{I}$ $(3)~\sum_{j=1}^{J}r_{mj}\;=\;\sum_{k=1}^{K_{m}^{\rm max}}k\,y_{mk}\quad\forall m\in\mathcal{I}$ $(4)~\sum_{m\in\mathcal{I}}r_{mj}\;\leq 1\quad\forall j$ $(5)~r_{mj_{1}}+r_{mj_{2}}-1\;\leq\;z_{mj_{1}j_{2}}\;\leq\;r_{mj_{1}},\;r_{mj_{2}}$ $\qquad\forall m\in\mathcal{I},~j_{1}<j_{2}$ $(6)~z_{mj_{1}j_{2}}+y_{mk}-1\;\leq\;x_{mkj_{1}j_{2}}\;\leq\;z_{mj_{1}j_{2}},\;y_{mk}$ $\qquad\forall m\in\mathcal{I},~k,~j_{1}<j_{2}$
Algorithm 4 Single-QCirc Scheduler
1:Input: $\mathcal{QC}$, $\mathcal{QP}$, $G$
2:while not IsEmpty($\mathcal{QC}$do
3:  for each $\text{QCirc}_{m}$ in $\mathcal{QC}$ do
4:   AssignQCirc($\text{QCirc}_{m}$)
5:   RemoveAssigned($\text{QCirc}_{m}$)
6:  end for
7:end while
III-B Single-QCirc Scheduling
Another approach for scheduling quantum circuits is to schedule them individually rather than in batches. In this case, we propose a single‑Qcirc scheduling algorithm, presented in pseudocode in Algorithm 4. This algorithm uses an MILP‑based assignment method to optimise QPU allocation for each circuit $\rm{QCirc}_{m}$, taking into account the available QPUs, network topology, and link characteristics. The key difference from the batch‑Qcirc scheduling method described in the previous subsection is that the batch approach jointly optimises allocation for all circuits in a batch, whereas the single‑circuit method follows a one‑by‑one, online strategy that locally optimises each assignment. As a result, diversity in circuit structures and characteristics across the batch is not fully exploited in this method. On the other hand, the single‑QCirc approach offers lower computational complexity and enables immediate scheduling of each circuit as soon as the required resources become available.
We adopt the same MILP‑based framework to determine the optimal allocation for each circuit. This assignment procedure is referred to as AssignQCirc in Algorithm 4, and its MILP formulation is provided in Formulation III-B. The same objective function and constraints from the batch formulation are used, but simplified for the single‑circuit case. Note that AssignQCirc is also used within the FILLQPU procedure described in Algorithm 3.
@@ III-B Single-QCirc Scheduling
Formulation 2. MILP-based Single-QCirc assignment problem $\min~\bigg{\{}\sum_{1\leq j_{1}<j_{2}\leq J}z_{mj_{1}j_{2}}\big{(}\omega_{0}\frac{w_{m}T_{j_{1}j_{2}}}{T_{\rm dec}}+\omega_{1}(1-F_{j_{1}j_{2}})\big{)}\bigg{\}}$ s.t. $(1)~\sum_{j=1}^{J}r_{mj}s_{j}N_{j}\;\geq\;w_{m}\quad\forall m\in\mathcal{I}$ $(2)~\sum_{j=1}^{J}r_{mj}\;\leq\;K_{m}^{\rm max}\quad\forall m\in\mathcal{I}$ $(3)~r_{mj_{1}}+r_{mj_{2}}-1\;\leq\;z_{mj_{1}j_{2}}\;\leq\;r_{mj_{1}},\;r_{mj_{2}}$ $\qquad\forall m\in\mathcal{I},~j_{1}<j_{2}$
IV Evaluation and results
@@
This section evaluates the proposed scheduling algorithms through simulations. We consider a DQC network consisting of 16 QPUs within a data centre. The QPUs have the following capacities: four QPUs with a capacity of 8 qubits, four with a capacity of 12 qubits, four with a capacity of 16 qubits, and four with a capacity of 20 qubits. These capacities are randomly assigned to individual QPUs using a fixed random seed to ensure consistent QPU configurations across all simulations. The entire simulation process is repeated for 10 different seed values, and the average results are presented.
@@
The network adopts a fat tree topology of four pods with each pod consisting of 4 QPUs, as illustrated in Fig. 4(a). In Sec. II, we have defined two key parameters characterising the links in the logical topology: time required to establish entanglement ($T_{j_{1}j_{2}}$) and fidelity ($F_{j_{1}j_{2}}$) of entanglement generation for remote gates. Here, we specify their values for the 4-pod fat-tree topology used in our simulations. This topology features three types of links, distinguished by the number of optical switches along the path—specifically, 1, 3, or 5 switches. We denote the corresponding latency values by $T_{\text{link}_{1}}$, $T_{\text{link}_{3}}$, and $T_{\text{link}_{5}}$, respectively. Assuming an elementary link latency of $T_{\text{el}}$ (corresponding to entanglement generation time between QPUs without loss from any optical switches), and noting that the entanglement generation rate is directly affected by link loss, thus the latency for each link type is calculated using the formula:
$T_{\text{link}_{n_{s}}}=\frac{T_{\text{el}}}{\eta_{s}^{n_{s}}},$ (7)
@@
where $\eta_{s}$ is the switch transmission efficiency and $n_{s}\in\{1,3,5\}$ is the number of switches in the path. In our simulations, we assume that all optical switches are of the same type and have an identical loss of $0.5$ dB. Given the short physical distances between QPUs in a data centre environment, propagation delays are considered negligible compared to the time needed for establishing entanglement. For the fidelity parameter, we assume a slight fidelity degradation with increasing link distance and loss, as follows: $F_{\text{link}_{1}}=0.96$, $F_{\text{link}_{3}}=0.94$, and $F_{\text{link}_{5}}=0.92$.
Figure 6: Connectivity graphs for benchmark circuits ($w=4$). Edge weights indicate the number of two‑qubit gate occurrences.
To establish the set of quantum circuits $\mathcal{QC}$, we use benchmark circuits from the Munich Quantum Toolkit [32]: Quantum Fourier Transform (QFT), Deutsch–Jozsa (DJ), W-state, and GHZ-state circuits. Fig. 6 shows connectivity-graph representations for $w=4$ qubits. From this figure, it can be inferred that GHZ and W-state circuits are sparsely connected, QFT circuits, modeled with a complete graph, are highly connected, and DJ circuits exhibit an intermediate level of connectivity. A set $\mathcal{QC}$ of size $M$ includes all four circuit types, with each type contributing $25\%$ of the total. For circuit ${\rm QCirc}_{m}$, the number of qubits $w_{m}$ is randomly selected from a type-specific range. Two types of required qubit range are considered. In Scenario 1 (Sc.1), the ranges of required qubit are $R_{\text{GHZ}}=R_{\text{WState}}=[18,26]$, $R_{\text{DJ}}=[14,22]$, and $R_{\text{QFT}}=[10,18]$, where smaller ranges are assigned to QFT and DJ to account for their higher connectivity requirements. In Scenario 2 (Sc.2), increased qubit demand is considered by shifting each range upward by 4, giving $R_{\text{GHZ}}=R_{\text{WState}}=[22,30]$, $R_{\text{DJ}}=[18,26]$, and $R_{\text{QFT}}=[14,22]$. For a given $M$, simulations are repeated over 10 independently generated $\mathcal{QC}$ sets, and results are averaged across runs.
TABLE II: Linear regression results for our benchmark circuits and simulation scenarios.
$\hat{\chi}_{0,k}$ $\hat{\chi}_{1,k}$ $\hat{\chi}_{2,k}$ $\hat{\chi}_{3,k}$ $R^{2}$ score
$k=2$ 0.0272 0.4345 0.0163 0.0434 0.9965
$k=3$ 0.1185 0.4808 0.0534 0.0802 0.9929
$k=4$ 0.1887 0.4585 0.081 0.1235 0.9845
The linear regression model is implemented using the scikit-learn library and a dataset covering all four circuit types with $w_{m}\in[10,30]$. The cut size for $k$-way partitioning, ${\rm Cut}_{m}^{(k)}$ in (4), is obtained by iteratively applying the K-L algorithm in NetworkX [33] until the desired number of partitions are obtained. The partitioning cost $\nu_{mk}$ is computed using the linear model in (5), with connectivity features $\gamma_{m}$, $\lambda_{2,m}$, and $\sigma_{m}/g_{m}$ extracted from each circuit’s graph. The regression coefficients and $R^{2}$ scores are reported in Table II, which indicate good predictive performance. Among the features, $\lambda_{2,m}$ is the most influential, while the coefficients for $\gamma_{m}$ and $\sigma_{m}/g_{m}$ are smaller but tend to increase with higher partition counts $k$.
The proposed scheduling methods (Algorithms 1 and 4) are implemented in Python. The parameter $\beta$ in the batch selection process of Algorithm 1 is set to $0.85$, and the partition count threshold $K^{\text{max}}_{m}$ is fixed at $4$ for all circuits. The MILP optimisation problems defined in Problem Formulations 1 and 2 are solved using the Python-MIP package with the Gurobi optimiser. After obtaining the optimal assignment, each quantum circuit is partitioned individually using the K–L algorithm, which determines both the number of remote gates between each QPU pair allocated to a given circuit and the total number of remote gates for that circuit.
Figure 7: Average number of remote gates in DQC divided by the total number of QCircs $M$, for the proposed and baseline scheduling methods. (a) Qubit-range scheme Sc.1, (b) Qubit-range scheme Sc.2.
Figure 8: Average number of circuit partitions divided by the total number of QCircs $M$, for the proposed and baseline scheduling methods. (a) Qubit-range scheme Sc.1, (b) Qubit-range scheme Sc.2.
Figure 9: Average Job Execution Time (JET) (normalised by $T_{\text{dec}}$) for the proposed and baseline scheduling methods across different circuit types. (a)–(d) Qubit-range scheme Sc.1, (e)–(h): Qubit-range scheme Sc.2.
@@
To evaluate the performance of the proposed scheduling methods, we consider several key figures of merit. All results are averaged over repeated simulation runs. We choose random scheduling as the baseline for comparison.
@@
[…]
  • Number of remote gates for each QCirc type: Average number of remote gates for a given QCirc type, computed over all its occurrences within $\mathcal{QC}$.
@@
  • Normalised Throughput: defined as the ratio of the number of QCircs ($M$) to Normalised Makespan. This indicates how many QCircs can be executed per unit time.
The normalised JET for a specific QCirc is obtained using the Qiskit framework [34], which represents the circuit as a sequence of layers, where each layer comprises quantum gates that can be executed in parallel. The execution time is calculated under the following conditions. First, we assume that remote gates within a layer of a given QCirc are executed sequentially rather than in parallel. Second, all local gates are assumed to have an identical execution duration. Based on these, we iterate through all circuit layers and identify whether each layer contains any remote gates or only local gates. Let $N_{\rm LL}$ denote the number of layers containing only local gates. We compute the normalised JET using the following expression:
@@
Fig. 7 and 8 present the average number of remote gates and circuit partitions per QCirc, respectively, for the proposed scheduling methods, compared against random scheduling as the baseline. The results are shown for qubit-range schemes Sc.1 and Sc.2 used in our simulations. For the Batch-QCirc scheduling method, several values of the parameter $\alpha$ (fraction threshold of the total qubits across all QPUs available before executing the next batch) are considered. As shown, both the Batch-QCirc and Single-QCirc methods result in a lower number of remote gates and, consequently, require fewer partitions for each QCirc compared to the baseline. Moreover, in all cases, the Batch-QCirc method outperforms the Single-QCirc approach with a more complex MILP optimisation process. The number of remote gates directly reflects the demand for communication resources between QPUs. Therefore, the results demonstrate that both proposed methods, and especially the Batch-QCirc assignment, effectively reduce inter-QPU communication overhead.
@@
Figure 10: Average makespan (normalised by $T_{\text{dec}}$) for the proposed and baseline scheduling methods. (a) Qubit–range scheme Sc.1, (b) Qubit–range scheme Sc.2.
Figure 11: Average throughput (normalised by $T_{\text{dec}}$) for the proposed and baseline scheduling methods. (a) Qubit-range scheme Sc.1, (b) Qubit-range scheme Sc.2.
Figure 12: Average makespan (normalised by $T_{\text{dec}}$) for Single-QCirc and Batch-QCirc ($\alpha=0.55$) scheduling methods under varying optical switch loss, assuming the qubit-range scheme Sc.2.
In our simulations, we also obtained the number of remote gates for all four evaluated circuit types individually within the set of circuits $\mathcal{QC}$. The overall results followed a similar pattern to those of JET (presented in Fig. 9)), as it also can be expected from Eq. (8). Therefore, for the sake of brevity, we do not include the results for all cases here, but present the results for the specific case with the total number of QCircs ($M$) equal to $36$ in Table III. For QFT when the total number of qubits required is low (in the range of 10–18 in Sc.1), the proposed Batch-QCirc assignment typically assigns QFT to a single QPU rather than distributing it across multiple QPUs, resulting in an average of zero remote gates. When the qubit demand is increased in Sc.2, however, the number of remote gates rises significantly across all scheduling methods.
TABLE III: Average number of remote gates for each QCirc type under the two qubit-range schemes Sc.1 and Sc.2.
Baseline Single-QCirc Batch-QCirc ($\alpha=0.55$) Batch-QCirc ($\alpha=0.65$) Batch-QCirc ($\alpha=0.75$)
Sc.1 Sc.2 Sc.1 Sc.2 Sc.1 Sc.2 Sc.1 Sc.2 Sc.1 Sc.2
QFT 106.3 199.4 4.37 24.76 0 16.65 0 16.1 0 17.3
DJ 9.6 13.74 1.92 4.53 1.36 4.41 1.17 4.45 1.25 4.3
W-State 2.37 2.83 1.56 2.18 1.5 2.08 1.17 2.08 1.25 2.09
GHZ 1.18 1.39 0.76 1.07 0.81 1.08 0.8 1.06 0.8 1.04
Next, we evaluate the normalised Job Execution Time (JET), obtained for each circuit type individually. As shown in Fig. 9, the proposed methods consistently outperform the baseline with reduced execution times across all QCirc types. Batch-QCirc scheduling further improves JET in most cases compared to the Single-QCirc approach. For Sc.1, Batch-QCirc significantly reduces the normalised JET of QFT (to about $0.0145$) by executing it locally within a single QPU. With higher qubit demand in Sc.2, QFT (Fig. 9(e)) and DJ (Fig. 9(f)) exhibit higher JET than W-state (Fig. 9(g)) and GHZ (Fig. 9(h)), reflecting their greater connectivity, which induces more remote gates and hence longer execution times. In joint-optimisation-based Batch-QCirc scheduling, reducing the JET of highly connected circuits tends to prioritise QPUs with greater capacity, improving fidelity and reducing communication links, but limiting optimisation for less connected circuits such as GHZ. This trade-off is evident in both Sc.1 and Sc.2, where Single-QCirc sometimes yields lower JET for GHZ. Such interdependence highlights the system-level nature of Batch-QCirc scheduling, where resource trade-offs naturally arise.
[…]
Next, we examine the normalised makespan and throughput, key metrics for efficient scheduling and resource management, with results shown in Fig. 10 and Fig. 11. The proposed methods significantly improve both metrics, with the Batch-QCirc approach outperforming the Single-QCirc method. This demonstrates the system-level effectiveness of Batch-QCirc scheduling: reductions in JET due to fewer remote gates translate directly into shorter makespan and higher throughput.
@@
We also compare Batch-QCirc scheduling with varying values of $\alpha$, examining all the results presented in this section. In terms of makespan and throughput, the Batch-QCirc scheduling with $\alpha=0.75$ consistently results in a slightly worse performance compared to other values. Since the parameter $\alpha$ determines when the algorithm proceeds to the next scheduling cycle, it plays a crucial role in affecting these evaluated metrics. For the results of average number of remote gates and JET, similarly, $\alpha=0.75$ does not show meaningful improvements over $\alpha=0.55$ or $\alpha=0.65$; In fact, the latter two often outperform $\alpha=0.75$. These results indicate that increasing waiting time for scheduling the next batch with a higher $\alpha$ may lead to unnecessary resource waste without performance gains.
Lastly, to investigate the impact of optical switch on the performance of the proposed DQC scheme, we increase the optical switch loss, which consequently raises the average entanglement generation time - $T_{\text{link}_{1}},T_{\text{link}_{3}},T_{\text{link}_{5}}$. Fig. 12 presents the makespan results for the single-QCirc and batch-QCirc ($\alpha=0.55$) methods under varying optical switch losses, $\eta_{s}$ = 0.5 dB, 1 dB, and 2 dB, assuming the qubit-range scheme Sc.2. The results demonstrate that as $\eta_{s}$ increases, the relative advantage of Batch-QCirc becomes more pronounced. For instance, for $M=36$, Batch-QCirc reduces the makespan by $14.7\%$ compared to Single-QCirc at $\eta_{s}=0.5~\mathrm{dB}$, with the improvement increasing to $28.3\%$ and $43.8\%$ at $\eta_{s}=1~\mathrm{dB}$ and $\eta_{s}=2~\mathrm{dB}$, respectively. These results highlight the particular promise of the Batch-QCirc approach in topologies with heterogeneous link qualities.
V Conclusion